# PrivacyRestore: Privacy-Preserving Inference in Large Language Models via Privacy Removal and Restoration

### v3

@@ C.2 Activation Steering Technique
where $\mathcal{R}$ is the steering/restoration meta vector.
Appendix D Formal Definitions of CDP and LDP
D.1 Centralized Differential Privacy (CDP)
CDP protects individual privacy only after data has been aggregated in a central repository, which is defined as,
Definition D.1.
(CDP). A randomized mechanism $\mathcal{M}:\mathcal{Q}\to\mathcal{G}$ fulfills $(\epsilon,\delta)$-differential privacy if for all adjacent queries $Q,Q^{\prime}\in\mathcal{Q}$ and all possible query results $G\subset\mathcal{G}$,
$\mathbb{P}\left(\mathcal{M}(Q)\in G\right)\leq\exp(\epsilon)\mathbb{P}\left(\mathcal{M}(Q^{\prime})\in G\right)+\delta.$
CDP (Dwork et al., 2016) ensures that adversaries cannot distinguish between $Q$ and $Q^{\prime}$ based on $G$ due to the similar probabilities, meaning the query results are probabilistically indistinguishable. This prevents adversaries from inferring characteristics about the repository based on multiple queries and their results.
D.2 Local Differential Privacy (LDP)
However, privacy risks can also emerge during the data collection process itself, as attackers may intercept user inputs while they are being transmitted to the central repository. LDP (Duchi et al., 2013) protect user inputs during transmission process by ensuring that attackers cannot distinguish between any two adjacent inputs, which is defined as,
Definition D.2.
(LDP). A randomized mechanism $\mathcal{M}:\mathcal{I}\to\mathcal{O}$ fulfills $(\epsilon,\delta)$-differential privacy if for all adjacent inputs $I,I^{\prime}\in\mathcal{I}$ and all possible outputs $O\subset\mathcal{O}$,
$\mathbb{P}\left(\mathcal{M}(I)\in O\right)\leq\exp(\epsilon)\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right)+\delta.$
The mechanism $\mathcal{A}$ processes the user input before transmitting it. LDP ensures that, even if attackers intercept $O$, they cannot distinguish between the initial user input $I$ and the adjacent one $I^{\prime}$.
Appendix E Proof of Theorem 5.1
Proof of $d_{\chi}$-privacy. As shown in Definition 2.1, if the input length is 1, indicating a single token $i_{1}$, $d_{\chi}$ can be:
$\mathbb{P}\left(\mathcal{M}(i_{1})\in o_{1}\right)\leq\exp(\epsilon d_{\chi}(i_{1},i_{1}^{\prime}))\mathbb{P}\left(\mathcal{M}(i_{1}^{\prime})\in o_{1}\right),$
where $o_{1}$ is the possible output set for $\mathcal{M}(i_{1})$ and the privacy budget is $\epsilon d_{\chi}(i_{1},i_{1}^{\prime})$. When the input becomes the sequential tokens $I=\{i_{1},i_{2},...,i_{n}\}$ with corresponding output sets $O=\{o_{1},...,o_{n}\}$, the LCP for the sequence of length $n$ is:
$\mathbb{P}$ $\left(\mathcal{M}(I)\in O\right)$
$=\mathbb{P}\left(\mathcal{M}(i_{1})\in o_{1}\right)\cdot\mathbb{P}\left(\mathcal{M}(i_{2})\in o_{2}\right)\cdot$
$\qquad\qquad...\cdot\mathbb{P}\left(\mathcal{M}(i_{n})\in o_{n}\right)$
$\leq[\exp(\epsilon d_{\chi}(i_{1},i_{1}^{\prime}))\mathbb{P}\left(\mathcal{M}(i_{1}^{\prime})\in o_{1}\right)]\cdot$
$\qquad\qquad...\cdot[\exp(\epsilon d_{\chi}(i_{n},i_{n}^{\prime}))\mathbb{P}\left(\mathcal{M}(i_{n}^{\prime})\in o_{n}\right)]$
$=\exp\Big{[}\epsilon\big{(}d_{\chi}(i_{1},i_{1}^{\prime})+...+d_{\chi}(i_{n},i_{n}^{\prime})\big{)}\Big{]}\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right)$
$=\exp\Big{[}\epsilon\sum_{j=0}^{n}d_{\chi}(i_{j},i_{j}^{\prime})\Big{]}\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right),$
where the privacy budget is $\epsilon\sum_{j=1}^{n}d_{\chi}(i_{j},i_{j}^{\prime})$. Commonly, we use the Euclidean distance as the $d_{\chi}$ function and obviously $\sum_{j=1}^{n}d_{\chi}(i_{j},i_{j}^{\prime})\propto n$. Therefore, the privacy budget of LDP grows linearly with the length $n$.
Proof of LDP. As stated in Definition D.2, if the length of input is 1, corresponding to a single token $i_{1}$, LDP can be expressed as:
$\mathbb{P}\left(\mathcal{M}(i_{1})\in o_{1}\right)\leq\exp(\epsilon)\mathbb{P}\left(\mathcal{M}(i_{1}^{\prime})\in o_{1}\right)+\delta,$
where $o_{1}$ is the possible output set of $\mathcal{M}(i_{1})$ and the privacy budget is controlled by $(\epsilon,\delta)$. Considering the sequence tokens $I=\{i_{1},i_{2},...,i_{n}\}$ and corresponding output sets $O=\{o_{1},o_{2},...,o_{n}\}$, the CDP for the sequence of length $n$ can be written as:
$\mathbb{P}$ $\left(\mathcal{M}(I)\in O\right)$
$=\mathbb{P}\left(\mathcal{M}(i_{1})\in o_{1}\right)\cdot\mathbb{P}\left(\mathcal{M}(i_{2})\in o_{2}\right)\cdot$
$\qquad\qquad...\cdot\mathbb{P}\left(\mathcal{M}(i_{n})\in o_{n}\right)$
$\leq[\exp(\epsilon)\mathbb{P}\left(\mathcal{M}(i_{1}^{\prime})\in o_{1}\right)+\delta]\cdot$
$\qquad\qquad...\cdot[\exp(\epsilon)\mathbb{P}\left(\mathcal{M}(i_{n}^{\prime})\in o_{n}\right)+\delta]$
$=\exp(n\epsilon)\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right)$
$\qquad\qquad+\delta\cdot\sum_{i=1}^{n}\prod_{j\neq i}\mathbb{P}\left(\mathcal{M}(i_{j}^{\prime})\in o_{j}\right)+\delta^{2}\cdot\cdots,$
where $\delta$ is typically considered a very small value. When $\delta$ approaches 0, we consider only the first two terms and then,
$\mathbb{P}\left(\mathcal{M}(I)\in O\right)$ $\leq\exp(n\epsilon)\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right)$
$+\delta\cdot\sum_{i=1}^{n}\prod_{j!=i}\mathbb{P}\left(\mathcal{M}(i_{j}^{\prime})\in o_{j}\right),$
which indicating the privacy budget becomes $\big{(}n\epsilon,\delta\cdot\sum_{i=1}^{n}\prod_{j!=i}\mathbb{P}\left(\mathcal{M}(i_{j}^{\prime})\in o_{j}\right)\big{)}$, according to the Definition of LDP in Section D.2 . The first term $n\epsilon$ obviously grows linearly with the length $n$. The second term can be view as $\delta$ multiplied by $\sum_{i=1}^{n}\prod_{j!=i}\mathbb{P}\left(\mathcal{M}(i_{j}^{\prime})\in o_{j}\right)$. The second term summarizes $n$ multiplicative terms, each bounded within $(0,1)$. The second term can be approximately considered to grow linearly with $n$. Therefore, the privacy budget of CDP also grows linearly with the length $n$.
Proof of CDP. The definition of CDP, as shown in Section D.1, is similar to LDP, with the only difference being that CDP applies to the user query $Q=\{q_{1},q_{2},...,q_{n}\}$ rather than the text input $I=\{i_{1},i_{2},...,i_{n}\}$. If we consider the token $i_{n}$ as a sub-query $q_{n}$, then similarly the definition of CDP for sequential sub-queries $Q=\{q_{1},q_{2},...,q_{n}\}$ can be:
$\mathbb{P}\left(\mathcal{M}(Q)\in G\right)$ $\leq\exp(n\epsilon)\mathbb{P}\left(\mathcal{M}(Q^{\prime})\in G\right)$
$+\delta\cdot\sum_{i=1}^{n}\prod_{j!=i}\mathbb{P}\left(\mathcal{M}(q_{j}^{\prime})\in g_{j}\right),$
where $G=\{g_{1},g_{2},...,g_{n}\}$ are the possible output sets for sequence inputs $Q$. The privacy budget is $\big{(}n\epsilon,\delta\cdot\sum_{i=1}^{n}\prod_{j!=i}\mathbb{P}\left(\mathcal{M}(q_{j}^{\prime})\in g_{j}\right)\big{)}$ and also grows linearly with the length $n$.
Appendix F Proof of Theorem 5.2
As shown in Figure 1, during the inference stage, only the meta vector and the query with privacy spans removed are transmitted from the client to the server. The meta vector holds information about all privacy spans and could be vulnerable to interception by adversaries who may attempt to reverse-engineer these spans.
PrivacyRestore protects the meta vector by adding noise $\mathcal{N}$ which is sampling from the distribution $p(\mathcal{N})\propto\exp(-\epsilon\|\mathcal{N}\|)$, before transmission, as shown in Eq 6. Firstly, to simply the question, we only consider the situation of only one edited head $h$. Assume $Z_{h}$ represents the normalized weighted sum of restoration vectors of all privacy spans on head $h$ without adding noise, $\mathcal{R}_{h}$ denotes the vector on head $h$ after adding noise, as shown in Eq 5 and 6. The process of adding noise can be represented by $\mathcal{M}$. Then, the possibility that $Z_{h}$ becomes $\mathcal{R}_{h}$ after adding noise $\mathcal{N}$ is
$\mathbb{P}(\mathcal{M}(Z_{h})=\mathcal{R}_{h})$ $=\mathbb{P}(Z_{h}+\mathcal{N}=\mathcal{R}_{h})$
$=\mathbb{P}(\mathcal{N}=\mathcal{R}_{h}-Z_{h})$
$=\exp(-\epsilon||\mathcal{R}_{h}-Z_{h}||).$
Then for any normalized weighted sums on head $h$, $Z_{h}$ and $Z^{\prime}_{h}$, we have
$\frac{\mathbb{P}[\mathcal{M}(Z_{h})=\mathcal{R}_{h}]}{\mathbb{P}[\mathcal{M}(Z^{\prime}_{h})=\mathcal{R}_{h}]}$ $=\frac{\exp(-\epsilon||\mathcal{R}_{h}-Z_{h}||)}{\exp(-\epsilon||\mathcal{R}_{h}-Z^{\prime}_{h}||)}$
$=\exp(\epsilon(||\mathcal{R}_{h}-Z^{\prime}_{h}||-||\mathcal{R}_{h}-Z_{h}||))$
$\leq\exp(\epsilon||Z^{\prime}_{h}-Z_{h}||).$
According to the definition of $d_{\chi}$-privacy in Section 2.1, the mechanism $\mathcal{M}$ satisfies $d_{\chi}$-privacy. In other words, by adding noise $\mathcal{N}$, adversaries cannot infer $Z_{h}$ from $\mathcal{R}_{h}$ even if $\mathcal{R}_{h}$ is intercepted. Moreover, the privacy budget of our methods is $\epsilon||Z^{\prime}_{h}-Z_{h}||$.
Then for multiple edited heads, the only difference is to concatenated all restoration vectors $\mathcal{R}_{h}$ to form a single vector, as shown in Section 4.2. Then, the concatenated vector is also added with noise to protect privacy and the privacy budget of our methods become $\epsilon||Z^{\prime}-Z||$, where $Z^{\prime}$ and $Z$ represent any pair of normalized weighted sums of restoration vectors concatenated across all edited heads. It is independent of the input length $n$ and depends on the hyperparameter $\epsilon$ the term $||Z^{\prime}-Z||$. Therefore PrivacyRestore fulfills $d_{\chi}$-privacy and provides a privacy budget $\epsilon||Z^{\prime}-Z||$ which is independent of the input length and inherently addresses the problem of the linear growth of privacy budget.
Appendix G Top-K Heads Selector Algorithm
@@
In the section, we present the detail implementation of Top-K Heads Selector, as shown in Algorithm 1. Firstly, we initialize an empty score list $L_{h}$ for each head. Secondly, each privacy span $s$ has its corresponding top-K heads set $\mathcal{H}_{k}^{s}$. For each head $h$ in $\mathcal{H}_{k}^{s}$, we append $\text{Score}(h,\mathcal{H}_{k}^{s})$ into its score list $L_{h}$. $\text{Score}(h,\mathcal{H}_{k}^{s})$ is defined as the rank of head $h$ among $\mathcal{H}_{k}^{s}$ in ascending order based on the accuracy of the probe associated with the head $h$ and privacy span $s$. Thirdly, we calculate the average value of each score list $L_{h}$ as the score of the corresponding head $h$. Finally, we sort all heads in the LLM by the scores and pick up top-K heads as the common top-K heads set $\mathcal{H}_{c}$.
Algorithm 1 Top-K Heads Selector
@@
Input: $\mathcal{S}$ is the set of privacy spans; $\mathcal{H}_{a}$ is the set of all heads; $\mathcal{H}_{k}^{s}$ is the top-K heads set of the privacy span $s$; $\text{Score}(h,\mathcal{H}_{k}^{s})$ return the rank of head $h$ among $\mathcal{H}_{k}^{s}$ in ascending order based on the accuracy of the probe associated with the head $h$ and privacy span $s$. The score of the head with lowest accuracy is $1$. The score of the head with highest accuracy is $K$.
@@
2:  for  $s\;$ in $\;\mathcal{S}$ do
3:     for $h\;$ in $\;\mathcal{H}_{k}^{s}$ do
4:        Append $\text{Score}(h,\mathcal{H}_{k}^{s})$ into $L_{h}$.

### v5

@@ C.2 Activation Steering Technique
where $\mathcal{R}$ is the meta vector.
Appendix D Selecting the Most Relevant Heads
In this section, we provide the implementation details of the probe technique (Alain and Bengio, 2016; Tenney et al., 2019; Belinkov, 2022) to identify the most relevant attention heads for each type of privacy span.
Let $I_{all}={I_{1},...,I_{m}}$ represent the user inputs in the training set, where $m$ is the size of the training set. For a given privacy span type $c$, let $Y_{c}={y_{1},...,y_{m}}$ represent the corresponding labels, where $y_{i}=1$ if and only if the input $I_{i}$ contains a privacy span of type $c$.
For each user input $I_{i}$, we record the hidden state of the last token on each attention head. We then train a binary classifier for each head, tailored to the privacy span type $c$, as the probe. The probe takes the hidden state of the last token as input and predicts whether the input contains the privacy span of type $c$. The probe is formulated as:
$\mathcal{F}^{c}_{h}(\textbf{u}_{h})=\sigma(\theta_{h}^{c}\cdot\textbf{u}_{h}),$ (13)
where $\mathcal{F}^{c}_{h}(\cdot)$ is the probe of privacy span type $c$ on head $h$, $\textbf{u}_{h}$ is the hidden state of the last token on head $h$, $\theta_{h}^{c}$ are the parameters of the probe, and $\sigma(\cdot)$ indicates the sigmoid function.
A probe $\mathcal{F}^{c}_{h}(\cdot)$ with higher accuracy indicates a stronger correlation between the head $h$ and the privacy span type $c$. Therefore, we only select the top $K$ attention heads, with the highest accuracies, as the most relevant heads for the privacy span type $c$.
Appendix E The Algorithm of Common Top-K Selector
In this section, we present the detailed implementation of Common Top-K Selector algorithm, as shown in Algorithm 1.
@@ Appendix E The Algorithm of Common Top-K Selector
Firstly, we initialize an empty score list $L_{h}$ for each head. Secondly, each privacy span type $c$ has its corresponding top-K heads set $\mathcal{H}_{k}^{c}$. For each head $h$ in $\mathcal{H}_{k}^{c}$, we append $\text{Score}(h,\mathcal{H}_{k}^{c})$ into its score list $L_{h}$. $\text{Score}(h,\mathcal{H}_{k}^{c})$ is defined as the rank of head $h$ among $\mathcal{H}_{k}^{c}$ in ascending order based on the accuracy of the probe associated with the head $h$ and privacy span type $c$. Thirdly, we calculate the average value of each score list $L_{h}$ as the score of the corresponding head $h$. Finally, we sort all heads in the LLM by the scores and pick up top-K heads as the common top-K head set $\mathcal{H}_{k}$.
Algorithm 1 Common Top-K Selector
@@
Input: $\mathcal{S}$ is the set of privacy spans; $\mathcal{H}_{a}$ is the set of all heads; $\mathcal{H}_{k}^{c}$ denotes the set of top-K heads corresponding to the privacy span type $c$; $\text{Score}(h,\mathcal{H}_{k}^{c})$ return the rank of head $h$ among $\mathcal{H}_{k}^{c}$ in ascending order based on the accuracy of the probe associated with the head $h$ and privacy span type $c$. The score of the head with lowest accuracy is $1$. The score of the head with highest accuracy is $K$.
@@
2:  for  $c\;$ in $\;\mathcal{C}$ do
3:     for $h\;$ in $\;\mathcal{H}_{k}^{c}$ do
4:        Append $\text{Score}(h,\mathcal{H}_{k}^{c})$ into $L_{h}$.